Subject: APO 3.5m users committee minutes, Dec 2013

From: Suzanne Hawley

Submitted: Thu, 19 Dec 2013 22:18:46 -0800 (PST)

Message number: 1324 (previous: 1323, next: 1325 up: Index)

PLEASE NOTE: The FITS headers for all instruments now use common time 
keywords, as discussed with the users committee and implemented during the 
2013 summer shutdown.  Below is a brief description, as documented on the 
3.5m current information page at:

We have recently had some questions about this change and will discuss it 
at the next users committee meeting.  Users are encouraged to provide 
feedback to their users committee members prior to the meeting (Jan 13, 


As of August 30, 2013, the 3.5-m has implemented a TAI time standard in 
all the instrument headers. All time keywords are referenced to TAI. In an 
effort to make all time keywords the same from instrument to instrument 
several keywords are no longer being written to some instrument headers. 
The standard keywords are: TIMESYS, UTC-TAI, UT1-TAI, LST, DATE-OBS (TAI 
time at start of exposure in year-month-day-hour:minute:second format), 
and EXPTIME. Every effort has been made to give users the minimum number 
of time related keywords with which they can recreate any of the old 
keywords on their own in post processing.


 		APO 3.5-m Users Committee Phonecon, 12/02/13

Attending:  Suzanne Hawley, John Bally, Rene Walterbos, Bill Ketzeback, 
Mary Beth Kaiser, Joe Huenerhoff, John Wisniewski, Joanne Hughes Clark, 
Scott Anderson, Michael Strauss, Jeff Bary, and Bruce Gillespie.

Absent:  Remy Indebetouw, Al Harper


User feedback and comments from institutional representatives:

o Washington (Scott Anderson) - Scott had nothing new to report.

o Chicago (Al Harper) - no report.

o New Mexico State (Rene Walterbos) - Rene had nothing new to report.

o Princeton (Michael Strauss) - Michael said that everything was fine.

o Colorado (John Bally) - John had nothing new to report.

o Virginia (Remy Indebetouw) - no report.

o Johns Hopkins (Mary Beth Kaiser) - Mary Beth had no concerns to report. 
She noted with appreciation that Hal Weaver got some very good data on 
comet ISON.

o Georgia State (Mike Crenshaw) - no report [Mike and other GSU 
astronomers were observing at APO at the time]

o Seattle Univ. (Joanne Hughes Clark) - Joanne said her runs have been 
fine, and the staff quite helpful.

o NAPG (Jeff Bary) & U. Oklahoma (John Wisniewski) - Jeff and John said 
that so far they have only been involved in the selection and scheduling 
of their upcoming observing programs and so far, so good.


Discussion of telescope/instruments report:

In Mark's absence, Suzanne pointed the committee to his report (see 
below).  One highlight is that the camera upgrade for the DIS slit viewer 
is delayed due to the camera manufacturer.  Bill mentioned that the 
increase in dark current seen in SPIcam images was correlated with the 
changeout of the ion pump about a year ago, and he's working with the 
manufacturer of the pump to find a remedy.  Suzanne emphasized the 
importance for SPIcam users of monitoring their data.  Joanne asked to 
please have the Observing Specialists let the users know when a 
dark-current "event" occurs.  Bill said he would ensure that this happens.

Mark Klaene's detailed report for the reporting period follows:


 	      3.5-m Telescope and Instruments Highlights
 	                10/15/13 through 11/25/13

0) Overview

Generally a productive period with the first snowfall of the season now on 
the ground, albeit not much snow.

1) Telescope

Telescope has been working very well.  We had one instance of the glycol 
fluid in the moat seal flowing into the main azimuth bearing. It took 2 
days to clean and repack the bearing. One night would have been lost 
however weather was so bad that we would not have been able to open 

2) Instruments

DIS has been operating normally.  The movement of DCam (shown in the mask) 
was located and a temporary fix put in place. Delivery of the new DCAM 
camera has been delayed by the manufacturer by about an additional month.

Agile is operational with no problems

TripleSpec had no operational problems this period.

SPIcam variable dark current has been isolated to the current ion pump 
type and controller.  A different ion pump has been installed but so far 
only minimal improvement has been noted. We are continuing to monitor dark 
current and ion pump spikes closely to understand the issues and determine 
best path forward.

Echelle is operational with no problems

NIC-FPS is operating in shared-risk observing mode due to occasional image 
corruption issue.  NFS less than 9 is operational.


ARCTIC PDR report:

Suzanne reported that the preliminary design review for the ARCTIC visible 
imager instrument was held at the end of October.  The review went well, 
and the review committee was supportive of the overall design and 
implementation approach.  There is an unresolved question of whether to 
use an e2v CCD or one from a relatively new vendor, Semiconductor 
Technologies Associates (STA).  Joe Huenerhoff is looking into the 
trade-space comparisons between the two.  The JHU Instrument Development 
Group will be asked to do a followup check of the optical design. 
Suzanne asked if anyone has any experience with STA detectors, please let 
us know.  The final PDR report will be made available to the users 


ARCSAT update:

Joe announced that the web interface for ARCSAT (the .5-m telescope at 
APO) is up and running; he is working on some minor bugs in the focus and 
dome controls.  We plan engineering testing to continue this week, and 
then will release the telescope to some experienced users for further 
tests.  Suzanne added that we planto finish the user documentation and 
testing in the next 1-2 months, and then will release the telescope to the 
general ARC user community for shared-risk observing programs, with the 
goal of regular observing starting in Q2.


ARC Board meeting summary:

Rene commented that the annual ARC Board of Governors meeting went well, 
and that the Board is happy with the progress on operational funding and 
prospects for the next two to three years.  The Board spent time 
discussing future issues, and was pleased that the current institutional 
partnership is stable and that there is real interest in expanding the 
consortium.  The Board was also satisfied with the continued high level of 
observatory performance and staff support; the operating budgets for next 
year were unanimously approved.  Bill noted that this was the fourth year 
running in which the 3.5-m telescope experienced a downtime of <1%.  It 
was also noted that SDSS4 was given Board approval to begin its first year 
of observing operations in mid-2014.


Q1 schedule:

Suzanne said that the Q1 schedule will be out shortly.  The Chinese 
consortium has signed the agreement to use the telescope; their program is 
denoted by IH01 in the Q4 2013 observing schedule and in the upcoming Q1 
2014 schedule.

On-site visits, training:

Suzanne reminded the committee, and by extension, the general user 
community, that the 3-night on-site training policy is applicable to all 
users, including the new leasing partners.  With more new people coming to 
the site, it is even more essential to note on the observing proposals and 
schedules if the observers will be on site at APO, and/or are there for 
training.  Additionally, if the new observer is a grad student, it is the 
policy that he or she be accompanied by a PhD astronomer from the 
student's institution.  APO staff are not expected to train students how 
to observe.  This led to the following action item:

===> ACTION:  Users Committee members to check the Q1 schedule to confirm 
that their observing programs conform to the on-site training policy [see]

Jeff asked a more general question, what is the documentation that new 
users should be looking at?  Suzanne said that the training link, plus the 
telescope and instrumentation user documentation at the APO website.  The 
committee agreed that it would be good to have a document that provides 
guidance (website links, etc.) to new users.

Additionally, it was noted that APO welcomes largish student and other 
groups to visit the observatory for multi-night observing programs and 
training.  These events need to have their arrangements made separately 
both for scheduling with Russet and Suzanne, and with Mark Klaene on-site, 
both well in advance.  John Bally said that Colorado is planning a student 
visit to APO in early April.  John Wisniewski has a group coming for a 
week sometime in Q2.  Jeff Bary is expecting his users to come to the site 
in groups of at most four people, on a yet-TBD schedule.


New business:

Rene and Suzanne pointed out that the 20th anniversary of the dedication 
of the 3.5-meter telescope is coming up this spring.  Discussions have 
begun on when, where, and how to celebrate this event.  Options include 
having a fete at APO in the spring or summer, including a celebratory 
dinner, and possibly a science meeting.  We have chosen not to do this in 
conjunction with the transition from SDSS3 to SDSS4 to avoid dilution of 
the 3.5-m celebration.  This leads to the following action item:

===> ACTION:  Users Committee members to think about the context and 
themes for a 20th-anniversary celebration for the 3.5-m, to be discussed 
at the next Users Committee telecon.


[open from previous meeting]:

===> ACTION:  Find users to assist with instrument documentation.

 	Status:  Open.  Thanks to Michael Strauss and Mary Beth, who have 
sent useful comments.

===> ACTION:  Users committee members to query their communities for 
interest in and volunteers to help with setting up a 3.5m user wiki.

 	Status:  Open.  No feedback yet.

[new from this meeting]:

 	===> ACTION:  Users Committee members to check the Q1 schedule to 
confirm that their observing programs conform to the on-site training 
policy [see]

 	===> ACTION:  Users Committee members to think about the context 
and themes for a 20th-anniversary celebration for the 3.5-m, and discuss 
at the next Users Committee telecon.


Next meeting:  The next Users Committee phonecon will be on Monday 13 
January at 8:30 AM Pacific Time.  The agenda and other materials will be 
sent to the committee members during the preceding week.

APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO  This is message 1324 in the apo35-general archive. You can find
APO  the archive on
APO  To join/leave the list, send mail to
APO  To post a message, mail it to